Calculate the moles of calcium chloride (CaCl2) needed to react in order to produce 85.00 grams of calcium carbonate (CaCO3).
1
Expert's answer
2020-10-16T08:56:23-0400

The reaction between the mentioned substances can be shown as following:

CaCl2 + Na2CO3 = 2NaCl + CaCO3

First, the mass of CaCO3 should be converted into moles:

n = m/Mr

where n - number of moles, m - mass, Mr - molecular weight

n(CaCO3) = m(CaCO3) / Mr(CaCO3) = 85.00 g /100.09 g/mol = 0.85 mol

As can be seen from the reaction, 1 mole of CaCl2 reacts to produced 1 mole of CaCO3. As a result, 0.85 mole of CaCl2 will also react to produce 0.85 mole or 85.00 g of CaCO3.


Answer: 0.85 mol